how do you convert y=-1/2x+5 to standard form?
Found 2 solutions by usyim88hk, AnlytcPhil:
Answer by usyim88hk(158)

y = (-1/2)x+5
(1/2)x+y = 5
Standard form is that u have to put x and y on the same sides of the equation

Answer by AnlytcPhil(1806)

how do you convert y=-1/2x+5 to standard form?
=====================================================
The standard form of the equation of a line is
Ax + B = C where |A|,|B|, and |C| are relatively prime integers,
if possible, and where A is positive.
y = (-1/2)x + 5
Clear of fractions by multiplying through by LCD = 2
2·y = 2·(-1/2)x + 2·5
2y = -x + 10
Add x to both sides:
x + 2y = 10
That's it!
Edwin